### v2

# Remote Cooling of Spin-ensembles through a Spin-mechanical Hybrid Interface

Abstract
We present a protocol for the ground-state cooling of a tripartite hybrid quantum system, in which a macroscopic oscillator acts as a mediator between a single probe spin and a remote spin ensemble. In the presence of weak dispersive coupling between the spins and the oscillator, cooling of the oscillator and the ensemble spins can be achieved by exploiting the feedback from frequent measurements of the single probe spin. We explore the parameter regimes necessary to cool the ensemble, the oscillator, or both to their thermal ground states. This novel cooling protocol shows that, even with only weak dispersive coupling, energy transfer-like effects can be obtained by simply manipulating the probe spin. These results not only contribute to the development of a practical solution for cooling/polarizing large spin ensembles, but also provide a relatively simple means of tuning the dynamics of a hybrid system. The proposed cooling protocol thus has broader implications for advancing various quantum technology applications, such as macroscopic quantum state generation and remote sensing.
preprint: APS/123-QED
I Introduction
Translating the potential of quantum technology into real-world applications remains a significant challenge due to the inherent limitations of current experimental platforms [1, 2]. One promising avenue is to explore the hybrid system approach, which promises to fully exploit the unique capabilities of different systems [3, 4, 5, 6]. In particular, high-fidelity entanglement between different systems can be established by exploiting their common coupling to a transducer, thereby generating indirect interactions [7, 6]. Transducers in the forms of nanomechanical oscillators (NMOs) [8, 9], magnons [10, 11], superconducting qubits [12, 13], superconducting resonators [14], optical cavities [15], and optomechanical cavities [16] have been considered to mediate spin–spin, spin-photon, and photon-phonon interactions.
However, it has been mainly proposed to use transducers to implement entanglement gates between remote qubits [16, 9, 10], and few transducer protocols focus on heat extraction from quantum systems to achieve ground state cooling or state preparation. The development of efficient polarization methods is particularly critical for spin ensembles, as conventional techniques often rely on resonant energy exchange methods, such as the Hartman-Hahn transfer [17, 18]. Polarization based on energy transfer typically exploits the interaction between a single spin and all the spins of the ensemble. This single spin can be polarized repeatedly, thus realizing a continuous removal of entropy until the ensemble is finally polarized. The challenge is that if the ensemble spins have similar g-factors, the risk of them being trapped in dark states due to Hamiltonian symmetries is high, preventing complete polarization of the ensemble [18, 19].
In this work, we consider a tripartite hybrid system where an oscillator serves as a transducer for a single probe spin and a spin ensemble. By exploiting the measurement feedback from the frequently measured probe spin, efficient heat extraction from the ensemble can be realized, requiring only weak dispersive (off-resonant) coupling between the spins and the oscillator. The probe control sequence is shown schematically in Fig. 1(b). Achieving complete ensemble polarization by such probe projections can avoid the dark state problem associated with resonant energy transfer. Furthermore, the ability to operate under dispersive interaction is important for various systems whose original dynamics do not allow direct energy exchange [20, 21]. Finally, we show that the same control sequence on the probe can also be used to prepare macroscopic oscillators (e.g., NMO) into specific quantum states that can potentially be used for sensing and error correction applications.
The oscillator here serves as a simplified conceptual model for a common bus (transducer) connecting two distant quantum systems that are not directly or weakly coupled. One can consider a general scenario where the bus is a large spin network that can be effectively approximated as a bosonic system via the Holstein-Primakoff transformation, such as the spin wave excitations (magnons) [22, 11]. This is relevant when a single spin, such as nitrogen-vacancy (NV) centers in diamond [23, 24], is used to polarize or probe distant target spins beyond the distance criteria set by the sensing volume [25]. The target may be distant carbon nuclear spins within the diamond substrate [26, 27], or spins attached to an external protein/molecule [28]. In such cases, probe-target interactions must be mediated by a transducer, such as the spin network schematically shown in the inset of Fig. 1(a) [26, 27, 28, 10]. Despite tuning both the network-to-probe and network-to-target interactions to resonance, the significant inhomogeneities typically present within a spin network make it difficult to become polarized, rendering resonant probe-target energy transfer ineffective [28, 26, 27]. Here, we instantiate such a scenario by considering the oscillator as the fundamental mode of an NMO, specifically the vibration of a clamped cantilever, as schematically shown in Fig. 1(a) [20, 21].
The paper is organized as follows: In Sec. II, we outline the Hamiltonian of the considered hybrid system and discuss possible experimental implementations of the proposed protocol. In Sec. A, we describe the dynamics of the hybrid system and analyze how periodic inversion pulses on the probe spin can significantly modify the dynamics of the ensemble and the oscillator. In Sec. IV, we detail how repeated probe projections can bring the oscillator and the ensemble into their ground states. In particular, we study the parameters that allow cooling of the spin ensemble, the oscillator, or both. Additionally, we discuss how it is possible to prepare the oscillator into a complex quantum state using the same probe control sequence as the cooling protocol. Finally, Sec. V provides a summary and discusses possible directions for future research.
II System Hamiltonian and experimental realization
For simplicity, we assume (though not necessary) that the ensemble spins interact with the common bosonic mode (oscillator) uniformly, with a coupling strength denoted as $g$. In addition, the probe spin is assumed to have a distinct coupling strength with the oscillator, denoted as $g_{0}$. The hybrid system’s Hamiltonian is then formulated as follows:
$H=\omega a^{\dagger}a+\sum_{k=0}^{N}\omega_{k}S_{k,z}+\left(g_{0}S_{0,z}+g\sum_{k=1}^{N}S_{k,z}\right)\left(a+a^{\dagger}\right),$ (1)
where $S_{i,z}$ signifies the z-component of the spin-1/2 operator for the ith spin, oscillating at the Larmor frequency $\omega_{i}$. The oscillator’s annihilation and creation operators are denoted by $a$ and $a^{\dagger}$, with a frequency $\omega$.
Figure 1: Schematic representation of the experimental setup and control strategy. (a) Shown is a clamped nano/micromechanical cantilever carrying a single defect spin, which in this study is exemplified by an NV electronic spin. The interaction between the NV center spin and the cantilever motion is mediated by strain, while single qubit gates are achieved by microwave (MW) pulses. Initialization and readout processes are facilitated by optical pulses, with a photon detector used for readout. A magnetic tip at the edge of the cantilever creates a gradient field that allows interaction with a nearby spin ensemble, coupling it to the mechanical motion of the cantilever. In the inset, we schematically represent a network of interacting spins as a quantum bus. When the number of spins in the network is large, it can be modeled as a bosonic system (e.g., magnons) by the Holstein-Primakoff transformation. Note that it is simplified here as the fundamental mode of the cantilever’s vibration. (b) The control sequence of the probe spin to perform the cooling protocol. First, the probe spin is reset to the |$\ket{0}$ state by optical pulses, then a $\pi/2$ MW pulse brings it to a state of equal superposition, which is followed by $n$ inversion MW pulses applied periodically with a carefully chosen time interval $\tau$. Another $\pi/2$ MW pulse is applied before the probe is optically read into the computational base. The above process is repeated $M$ times, bringing the oscillator and the spin ensemble progressively closer to their thermal ground states. Note that the probe to be read must be post-selected in |$\ket{0}$ each time, otherwise the whole process starts over.
This hybrid system can be realized as shown in Fig. 1(a) by a unilaterally clamped cantilever, embedding a single NV center positioned at the clamping point such that it experiences maximum strain and thus higher coupling strength. On the free side of the cantilever, we attach a magnetic tip to generate a strong field gradient that couples to a nearby spin ensemble on an external substrate [9, 6]. In this way, we exploit both possibilities to achieve spin-mechanical coupling [7, 29, 30] by local (strain) and non-local (magnetic-tip) coupling of spins to mechanical modes within a single setup. This allows the distinct manipulation of quantum systems that are otherwise not controllable.
The probe spin is chosen as the electronic spin associated with a NV center due to its good spin coherence properties, high degree of spin control, and well resolved optical spectra at low temperatures [31, 32, 33]. Its optical interface allows excitation to different optically excited energy levels that can either initialize or read out the spin state [34]. The NV electronic spins are also sensitive to their environment, allowing them to be (weakly) coupled to external spins, photons [20, 35] and phonons [21, 16].
The NV electronic spin is spin-1, which can be operated as a qubit with its $m_{s}=0,-1$ sublevels. This requires an external magnetic field aligned with its symmetry axis to separate the $m_{s}=\pm 1$ levels, which is typically around 400 Gauss [32]. Such an arrangement negates the need to consider transverse spin-strain coupling due to its direct interaction between the $m_{s}=\pm 1$ levels [7, 29, 30], allowing us to safely ignore the $m_{s}=+1$ sublevel in this work. What remains is the longitudinal strain coupling acting on the qubit levels, with a number that can be up to 10 kHz for NV centers [36]. This number can be even higher for other solid-state defect centers [37].
In contrast, the spin ensemble to be cooled can be chosen more flexibly, since precise control of these spins is not a requirement for the current study. Potential targets include distant spins on other coupled cantilevers, spins at the clamped edge that are strain-coupled to the mechanical mode, or surface electron spins that are not well controlled. For example, if the ensemble spins are dark spins on the surface of the NMO [38] or other ferromagnetic particles [39], optical manipulation over them can be difficult, highlighting the need for other indirect polarization techniques. Note that this ensemble-oscillator interaction generated by the attached magnetic tip can be (1-2 orders of magnitude) stronger than the spin-strain coupling between the probe and the oscillator [40].
It is important to note that although we specifically consider a cantilever system, an alternative experimental realization could use the spin wave excitations of the spins (magnons) as the transducer [10, 22, 11]. An advantage of this alternative is that the spin-magnon coupling can easily reach a few MHz [41], much stronger than the typical strength of spin-cantilever coupling by either strain or magnetic tip.
III System dynamics and its tuning
A Original system dynamics
In the following, we will consider the system in the rotating frame of precession frequencies of the oscillator and the spins. The rotating-frame Hamiltonian reads
$\tilde{H}=\left(g_{0}S_{0,z}+g\sum_{k=1}^{N}S_{k,z}\right)\left(\hat{a}e^{-i\omega t}+\hat{a}^{\dagger}e^{i\omega t}\right).$ (2)
Since the ensemble spins form a finite-dimensional Hilbert space, it is convenient to write the system evolution operators in the form [42, 43]:
$\begin{split}U(t)&=\ket{0}\bra{0}\otimes\sum_{\scriptscriptstyle k=0}^{\scriptscriptstyle N}\mathcal{D}_{k,+}(t)\otimes\mathrm{I}_{k}\\ &\qquad\qquad+\ket{1}\bra{1}\otimes\sum_{\scriptscriptstyle n=0}^{\scriptscriptstyle N}\mathcal{D}_{n,-}(t)\otimes\mathrm{I}_{k},\\ &\mathcal{D}_{k,\pm}(t)=\mathcal{T}\exp\left(\pm i\int_{0}^{t}\mathrm{d}t^{\prime}\ h_{\scriptscriptstyle k,\pm}(t^{\prime})\right),\end{split}$ (3)
where $\mathcal{T}$ the time ordering operator; the terms $h_{\scriptscriptstyle k,\pm}(t)$ are the oscillator Hamiltonians conditioning on the spin states, which are expressed as:
$h_{\scriptscriptstyle k,\pm}(t)=\frac{g(k-N/2)\pm g_{0}}{2}\left(\hat{a}e^{-i\omega t}+\hat{a}^{\dagger}e^{i\omega t}\right);$ (4)
and the operator $\mathrm{I}_{\scriptscriptstyle k}$ projects the spin ensemble into the subspace where $k$ spins are pointing up. For example, when $N=3$ and $k=2$, the projector is written as
[…]
It has been shown that the Magnus expansion of the evolution operators can be simplified as [42, 43]
$\begin{split}\mathcal{D}_{\scriptscriptstyle k,\pm}(t&)=\exp\left(\pm i\int_{0}^{t}\mathrm{d}t^{\prime}h_{\scriptscriptstyle k,\pm}(t^{\prime})\right)\\ &\qquad\times\exp\left(\frac{1}{2}\int_{0}^{t}\mathrm{d}t^{\prime}\int_{0}^{t^{\prime}}\mathrm{d}t^{\prime\prime}\left[h_{\scriptscriptstyle k,\pm}(t^{\prime}),h_{\scriptscriptstyle k,\pm}(t^{\prime\prime})\right]\right)\\ &=\exp\left[\frac{g(k-N/2)\pm g_{0}}{2\omega}\left(\alpha(t)a^{\dagger}-\alpha^{*}(t)a\right)\right]\\ &\qquad\quad\times\exp\big{(}it\theta_{\scriptscriptstyle k}\big{)}.\end{split}$ (6)
where $\alpha(t)=1-e^{i\omega t}$ and $\theta_{\scriptscriptstyle k}$ is given by
$\begin{split}\theta_{\scriptscriptstyle k}&=-\frac{\omega t-sin(\omega t)}{\omega t}\frac{\left[g(k-N/2)\pm g_{0}\right]^{2}}{4\omega}.\end{split}$ (7)
B Inversion pulses on probe
In addition to the natural dynamics induced by the Hamiltonian given in Eq. (1), we also apply a periodic sequence of $\pi$ pulses to the probe spin. These pulses significantly alter the dynamics of the system by imposing a time-periodic function:
$f(t)=\begin{cases}+1&\text{if }2j\tau<t<(2j+1)\tau\\ -1&\text{if }(2j+1)\tau<t<(2j+2)\tau\end{cases}$ (8)
where $j=0,1,2,\cdots,n$, and $\tau$ denotes the interval between successive pulses. By replacing $g_{0}$ with $f(t)g_{0}$ in $h_{\scriptscriptstyle k,\pm}(t)$, the conditional evolution operators in Eq. (6) are now written as
$\begin{split}\mathcal{D}^{f}_{\scriptscriptstyle k,\pm}(t)=&\exp\Big{[}\frac{g(k-N/2)}{2\omega}\left(\alpha(t)a^{\dagger}-\alpha^{*}(t)a\right)\\ &\;\pm\frac{g_{0}}{2\omega}\left(F(\epsilon,t)a^{\dagger}-F^{*}(\epsilon,t)a\right)\Big{]}\times\exp\left[i\theta_{\scriptscriptstyle k}(t)\right],\end{split}$ (9)
where the filter function $F(t)$ is given by [42]
$\begin{split}F(t)&=i\omega\ \sum_{k=0}^{n}\int\limits_{k\tau}^{(k+1)\tau}\mathrm{d}t\;(-1)^{k}e^{-i\omega t}\\ &=\left[1-(-1)^{n+1}e^{-i(n+1)\omega\tau}+2\sum_{k=1}^{n}(-1)^{k}e^{-ik\omega\tau}\right].\end{split}$ (10)
This filter function is centered at $\tau=\pi/\omega$ with its width decreasing approximately as $1/n^{2}$ [43, 42, 44] and asymptotically converges to a delta function:
$\lim\limits_{n\rightarrow\infty}F(t)=2(n+1)\delta(\omega-\frac{\pi}{\tau}),$ (11)
which supports our initial assumption that all other phonon modes are effectively decoupled in Eq. (1) with large number of pulses $n$. In addition, we also want the maximum possible coupling strength between them. One can see from Eq. (9) that $|g_{0}F(t)|$ should be comparable to the oscillator frequency $\omega$, which requires at least $n\approx\mathcal{O}(\omega/g_{0}$) [43].
To achieve some tunability of the system dynamics that facilitates our proposed cooling protocol, we introduce the pulse detuning $\epsilon$ as a control parameter:
$\tau=\frac{\pi}{\omega-\epsilon},\text{ with }\frac{\epsilon}{\omega}\ll 1.$ (12)
Note that the detuning is intentionally kept significantly smaller than the oscillator frequency, so that the inversion pulses are in the near-resonance regime. Accordingly, the filter function can now be simplified as
$F(\epsilon,t=(n+1)\tau)=1-e^{i\epsilon t}+\frac{2(1-e^{i\epsilon t})}{e^{-i\epsilon\tau}-1}.$ (13)
For higher values of pulse detuning $\epsilon$, making the above equation closer to a delta function would require an even larger pulse number $n$.
Figure 2: (a) The quantum circuit progressively projects the oscillator ensemble system by projection of the probe spin. The pulse sequence on the physical level is shown in fig. 1(b). In each repetition, the probe spin is first prepared in the equal superposition state $\ket{+}$ by the optical reset and a microwave Hadamard gate. Then repeated inversion pulses are applied, resulting in a controlled oscillator displacement $\sum_{k}\mathcal{D}^{f}_{\scriptscriptstyle k,\pm}(t)\otimes\mathrm{I}_{k}$. Here the subscript $\pm$ depends on the spin state of the probe and $k$ on the number of upward-pointing ensemble spins. The procedure ends with a Hadamard gate and a measurement of the probe spin, with the result set to 0. Such a repetition effectively implements the projector given in Eq. (14). (b) An example of how an ensemble of $N=14$ spins is polarized after $M=100$ probe projections. The spins are initially in a completely mixed state, which is projected to the state with all spins pointing up with a probability of 95.82%. Note that achieving such a high degree of polarization requires a carefully chosen pulse detuning, which in this case is $\epsilon=1.16$ kHz. Further simulation details are summarized in Appendix A.
C Tuning dynamics through probe projections
Although we focus specifically on the weak dispersive coupling between each spin and the oscillator, we deliberately ignore any direct interactions between the spins, making energy transfer between the probe and the ensemble unlikely. Nevertheless, by carefully choosing the pulse detuning $\epsilon$ and the number of pulses $n$, energy transfer-like changes in the dynamics of the spin ensemble can be induced, as we will discuss below.
The dynamics modification is facilitated by a cyclic process, involving the projective measurement and reinitialization of the probe spin, as schematically shown in Fig. 1(a). First, the probe is prepared in the state $\ket{+}$. Then the system undergoes an evolution for a time $t=n\tau$ under repeated inversion pulses, as described in Eq (3). Finally, the probe spin is measured in the computational basis after an additional $\pi/2$ pulse is applied to it.
The process described above is represented by the circuit in Fig. 2(a), effectively implementing oscillator displacements conditioning on the probe state. The whole circuit is then equivalent to the implementation of a projector on the oscillator-ensemble system, which is mathematically formulated as:
$\mathcal{P}_{\scriptscriptstyle N,\pm}=\sum_{\scriptscriptstyle k=0}^{\scriptscriptstyle N}V_{k,\pm}\otimes\mathrm{I}_{k},\quad V_{k,\pm}=\frac{1}{2}\left(\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,+}(t)\pm\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,-}(t)\right).$ (14)
The operator $V_{k,\pm}$ is a projector on the oscillator, where $k$ is the number of ensemble spins pointing up, and the subscript $\pm$ depends on the measurement result of the probe spin.
Note that such repeated projections of the probe is central to the oscillator cooling protocol in our previous work [43], which is now extended to cooling/polarizing a spin ensemble. Here, we start with $N$ ensemble spins initially in a fully mixed state, and the oscillator initially in its thermal state. These initial states are written as
[…]
To cool the ensemble-oscillator subsystem, we repeatedly implement the circuit shown in Fig. 2(a). It is important to emphasize that, each time, only a probe measurement result of 0 is accepted, otherwise the whole process is restarted. This process effectively implements the projector $\mathcal{P}_{\scriptscriptstyle N,+}$ given in Eq. (14) many times. The probability that the ensemble is projected into the $\mathrm{I}_{k}$ sector after $M$ repetitions is calculated as
$\mathrm{P}_{m}(M)=\frac{1}{\sum_{m}\mathrm{P}_{m}}\mathrm{Tr}[\mathcal{P}_{\scriptscriptstyle N,+}^{M}\,\rho_{B}(0)\,\mathcal{P}_{\scriptscriptstyle N,+}^{M}],$ (16)
where $m=(2k-N)/2$ is the total magnetization of the ensemble. This dependence on $m$ shows that, with appropriate parameters, it is possible to polarize the ensemble, i.e., to obtain $\mathrm{P}{\scriptscriptstyle N/2}(M)\approx 1$ when $M$ is large enough.
IV System cooling
In this section, we describe how the repeated probe projections can impose a thermal filter on the hybrid system, allowing for controlled cooling of either the oscillator, the ensemble, or both. We focus primarily on the weak coupling regime, where the coupling strength $g$ is significantly smaller than the mechanical oscillation frequency $\omega$, denoted as $g\ll\omega$. To explore this regime, we set the spin coupling strengths $g,~{}g_{0}$ to $10$ kHz, while the mechanical oscillation frequency $\omega$ is chosen to be $1.2$ MHz. Although such a frequency for the mechanical oscillator is experimentally achievable [45, 46], the chosen values for the spin-mechanical couplings are at the upper end of what is typically achievable in practice [43].
These choices represent a compromise between stronger coupling for faster cooling and additional noise due to imperfect pulses on the probe. As we discussed in Eq. (11), at least $\mathcal{O}(\omega/g_{0})$ pulses are required for each probe projection to ensure that the effective spin-oscillator coupling is comparable to the oscillator frequency, which becomes clear in Eq. (9). However, a higher number of pulses would result in a longer total evolution time to implement the cooling protocol, which is undesirable in the presence of decoherence. Throughout this work, we set the number of pulses in each round of probe projection to $n=\omega/g_{0}=120$, as shown in Fig. 1(b).
Figure 3: Pulse Detuning: A key parameter in the cooling scheme. (a) This panel shows the intricate dependence of achieving complete spin polarization on pulse detuning ($\epsilon$). The probability that all ensemble spins are aligned upward after $M=100$ probe projections, denoted by $\mathrm{P}_{\scriptscriptstyle m=N/2}(M)$ and derived in Eq. (16), is plotted on the vertical axis. This probability is affected by both the number of spins in the ensemble and the size of $\epsilon$, with $\epsilon=2$ kHz already appearing too large for effective polarization of the ensemble. (b) If we choose an even larger detuning ($\epsilon=2.5$ kHz) and set the number of spins to $N=4$, we observe a rapid decrease of the thermal occupancy of the oscillator to zero, indicating that the oscillator is cooling. However, the spin polarization remains almost unchanged and reaches a steady state of 0.0625 after about 40 probe projections. Note that we have considered an ensemble of $N=4$ spins, so the y-axis represents the relative value of $\mathrm{P}_{z}(M)/2$. Correspondingly, the y-axis also represents the relative value of the thermal occupancy of the oscillator with respect to its initial value, which is set to about 45. Further simulation details are summarized in Appendix A.
A Pulse detuning as a control parameter
The probability $\mathrm{P}_{m}(M)$ in Eq. (16) depends non-trivially on the filter function in Eq. (13) that is imposed by pulses on the probe. As discussed in our previous work [43], varying the value of $\epsilon$ can lead to cooling, heating, or squeezing of the oscillator. In particular, Ref. [43] has found that a ratio of $\epsilon/\omega\approx 5\times 10^{-3}$ leads to rapid cooling of the oscillator, which corresponds to $\epsilon\approx 6$kHz for the oscillator frequency of $1.2$ MHz.
However, to achieve complete polarization of the ensemble, we must avoid premature cooling of the oscillator before the ensemble is polarized. This is due to the significant effect of the thermal occupancy of the oscillator on the indirect interaction strength between the probe and the ensemble, as elaborated in Eq. (1). In particular, at zero thermal occupancy, the oscillator-mediated effective interaction between the probe and the ensemble spins goes correspondingly to zero.
First, we run simulations with the pulse detuning $\epsilon$ close to zero to explore the near-resonance regime. By adjusting its magnitude from 0 to 2 kHz and varying the number of spins in the ensemble, our simulations reveal a nuanced relationship between the probability of achieving complete ensemble polarization and the pulse detuning $\epsilon$, as shown in Fig. 3(a). To cool the spin ensemble, the choice of $\epsilon$ should satisfy the condition that the probe-oscillator coupling is enhanced without prematurely cooling the oscillator. Consequently, satisfying this condition requires fine-tuning of all parameters of the Hamiltonian. This makes it difficult to determine the exact dependence of the optimal pulse detuning $\epsilon$ on the number of ensemble spins $N$ and the number of probe projections $M$, as shown in Fig. 3(a). For example, it can be seen that within an ensemble of constant size, varying the detuning can lead to very different results when implementing the proposed cooling protocol. Another important observation from Fig. 3(a) suggests that a detuning of $\epsilon=2$ kHz may be large enough to facilitate rapid oscillator cooling and thereby prevent ensemble polarization.
To investigate this further, we introduce an even larger detuning of $\epsilon=2.5$ kHz for an ensemble of $N=4$ spins. The simulation results shown in Fig. 3(b) indicate that at this substantial pulse detuning, the oscillator quickly reaches its ground state after about 80 probe projections. This premature cooling of the oscillator reduces the interaction between the probe and the ensemble, so that the ensemble remains largely unpolarized. The ensemble polarization fluctuates only slightly around the final saturation value for the first few tens of probe projections, as shown in the inset of Fig. 3(b).
Furthermore, as an example, we search for the optimal pulse detuning for cooling a larger ensemble of $N=14$ spins. An exhaustive parameter search leads us to set the detuning to $\epsilon=1.16$ kHz. This setting shows that the sector with all spins pointing up (i.e. $m=7$) is predominantly left after $M=100$ probe projections. The probability of full polarization is as high as 95.82%, as shown in Fig. 2(b).
The above results underscore the critical role of $\epsilon$ as a controlling parameter in the cooling protocol. In addition, a subtle and interesting aspect of cooling the oscillator is to choose the detuning $\epsilon$ precisely so that the conditional oscillator displacement operator is not perfectly aligned with the position or momentum quadrature. Otherwise, the oscillator will be squeezed instead of cooled [43]. Note that, in the resonant case that $\epsilon=0$, the displacement operator of the oscillator is along the position quadrature.
This observation is interestingly related to the encoding of Gottesman-Kitaev-Preskill (GKP) states, which have applications in quantum sensing and quantum error correction [47, 48, 49]. Its encoding can be realized by the same control of the probe spin as our cooling protocol. Starting with an oscillator in its zero photon ground state, repeated probe projections can actually increase the number of photons and leave the oscillator in GKP states [47, 48]. The generation of these states in macroscopic objects could enrich our understanding of the quantum-classical interface, notwithstanding their important roles in quantum error correction [47, 48] and quantum sensing [50]. More details of GKP encoding can be found in Appendix C.
Figure 4: Cooling of both the oscillator and the ensemble. This figure shows the polarization $\mathrm{P}_{z}(M)$ and the thermal occupancy $n(\omega)$ for both the spin ensemble and the oscillator, illustrating how the cooling/polarization proceeds with the number of probe projections $M$. The effective cooling of these quantum systems occurs in different parameter regimes. Initially, the spin ensemble is cooled by adjusting the pulse intervals to near-resonance conditions, specifically $\tau_{1}$ such that $\epsilon_{1}=1.1$ kHz. In this case, the thermal occupancy of the oscillator remains largely unaffected. After polarizing the spin ensemble, we change the pulse interval to $\tau_{2}$ with a pulse detuning of $\epsilon_{2}=5$ kHz. This more significant detuning leads to a rapid cooling of the oscillator to its ground state. Note that the number of ensemble spins is $N=4$ so that the y-axis represents the relative value of spin polarization $\mathrm{P}_{z}(M)/2$. The oscillator thermal occupancy is also plotted as its relative to the initial occupancy $n_{0}\approx 45$. Further simulation details are summarized in Appendix A.
B Simultaneous cooling of ensemble and oscillator
In this section, we discuss the cooling of both the spin ensemble and the oscillator simultaneously. The key to achieving this goal is to avoid premature cooling of the oscillator, which should instead occur after the ensemble is fully polarized. This requires tuning the pulse detuning $\epsilon$ separately for the cooling phases of the oscillator and the ensemble. Fig. 4 illustrates a pulse sequence that effectively polarizes an ensemble of $N=4$ spins while cooling an oscillator from an initial thermal occupancy of $n_{0}\approx 45$.
During this simulation, we find that the optimal detuning that achieves complete polarization of the ensemble is identified as $\epsilon_{1}=1.1$kHz, while the oscillator occupancy is not much reduced. In this case, performing 250 probe projections can achieve complete polarization of the ensemble. Then, to cool the oscillator, we adjust the detuning to a higher value, $\epsilon_{2}=5$kHz, and execute an additional series of $50$ probe projections. This transition in the magnitude of the pulse detuning results in a rapid cooling of the oscillator, as shown in the right part of Fig. 4.
It is important to note that increasing $\epsilon_{2}$ requires a higher number of inversion pulses on the probe to achieve a comparable cooling effect, as can be seen from Eq. (13). However, implementing a higher number of inversion pulses presents experimental challenges, primarily due to the potential for pulse imperfections. This consideration emphasizes the need for precise control and optimization in the experimental setup.
V Discussion
In this study, we have developed a novel ground-state cooling protocol for a tripartite hybrid quantum system consisting of a spin ensemble, an oscillator, and a probe spin. Central to this protocol is the exploitation of the feedback provided by frequent measurements of the prob spin. Our analysis has shown that even with weak dispersive coupling between the spins and the oscillator, one can effectively cool both the ensemble and the oscillator to their ground states, either separately or simultaneously. We would also like to emphasize that, except for the probe, there is no need for quantum coherence in other parts of the hybrid system, marking a clear distinction from resonant transfer schemes. The polarization considered here can also be viewed as a remote sensing scheme when considering a macroscopic oscillator, where the boundaries of the sensing volume can be extended beyond the limits set by phase accumulation schemes [25].
Our numerical simulations, performed over a range of parameters, underscore the feasibility of the cooling protocol, and lay the groundwork for its experimental realization. For example, our results indicate that a spin ensemble consisting of $N=14$ spins could be cooled with a spin-mechanical coupling strength of only a few kHz at a mechanical oscillation frequency close to 1 MHz. Although such coupling strengths are within the experimental reach, the cooling efficiency can be further improved by miniaturization of the mechanical oscillator [51]. In the context of solid-state defects in silicon carbide, coupling strengths can reach a few MHz [37], promising even more efficient cooling rates. In addition, the use of magnons instead of NMOs could significantly enhance the spin-oscillator coupling, facilitating the experimental adoption of our protocol.
A notable challenge of our cooling approach is its dependence on post-selection, which results in a decreasing success rate with increasing number of probe projections. Nevertheless, our protocol offers a significant advantage for cooling large spin ensembles that are traditionally difficult to handle and cannot be cooled by standard optical or energy transfer techniques. For spin ensembles that are amenable to conventional cooling methods, such as NV centers in diamond, our protocol adds to the cooling toolbox.
This research highlights the significant impact that precise control of a single spin can have on the dynamics of a hybrid system, enabling energy transfer-like effects beyond the inherent capabilities of weak dispersive coupling. In addition to remote sensing, these results also provide insight into the preparation of complex quantum states such as the GKP states in a macroscopic object. Furthermore, the ability to interact/communicate with a few target spins within a large network may also be useful in quantum network protocols where a similar result is desired.
Acknowledgement
The authors would like to thank Jixing Zhang, Xin Zhang, and Ruoming Peng for their valuable feedback on the manuscript. This work was supported by the German Federal Ministry of Education and Research (BMBF) through the project QECHQS with Grant No. 16KIS1590K, and through the Future Cluster QSens. Additional support was provided by the Baden-Württemberg Stiftung through the project SPOC with Grant agreement No. QT-6. The Deutsche Forschungsgemeinschaft (DFG) has funded this work under GRK 2642, and under Project No. 384846402 as part of FOR 2724.
Appendix A Simulation details
The coupling parameters are also $g_{0}=g=10$ kHz, with an oscillator frequency of $\omega=1.2$ MHz. Accordingly, we set the number of inversion pulses for each round of probe projection to $n=120$, see the probe control sequence in Fig. 11(b). This is to ensure that the effective probe-oscillator coupling is comparable to the oscillator frequency, as discussed in Eq. (11).
In this work, we perform the simulation by calculating the dynamics directly on the initial Hamiltonian given in Eq. (1) without any approximation. In addition, for simplicity, we approximate the oscillator mode as having only 100 levels, which is expected to be sufficient since we assume an initial oscillator thermal occupation of $n_{0}\approx 45$ in the simulations.
Appendix B Effect of decoherence
Our numerical simulations have not considered the direct effect of spin dephasing noise. For the hybrid system considered in this work, such noise is induced by the spin bath surrounding the probe and the ensemble, and is expected to be the dominant noise source for the hybrid system considered [52]. In each measurement, both the ensemble spins and the spin bath surrounding them collapse partially and simultaneously towards the steady states. This collapse is sensitive only to the probe population in the measurement basis [53].
Similarly, in the case of the oscillator, its dissipation ($T_{1}$) also has a negligible effect. As shown in Fig. 4, the cooling of the spin ensemble requires more measurements and time compared to the cooling of the oscillator. Since the (initial) state of the oscillator is unaffected during the cooling of the ensemble, the effect of the oscillator’s $T_{1}$ can be ignored. Counterintuitively, in some cases rapid decoherence of the intra-bath correlations can actually increase the cooling rate rather than decrease it [54]. See further discussion of the effect of oscillator $T_{1}$ decoherence (relaxation) on its cooling in our previous work [43].
Appendix C Encoding a qubit into a macroscopic oscillator
In this section, we consider a hybrid system consisting only of the probe spin and the oscillator. We will show that we can prepare a macroscopic mechanical oscillator into a complicated quantum state using the same probe control sequence as our cooling protocol in the main text. The exploration of such quantum superposition states in macroscopic objects has emerged as a significant area of interest, mainly because of its potential to explore the frontier between classical and quantum physics [55, 56].
The quantum state in which we propose to prepare into a macroscopic mechanical oscillator is the so-called Gottesman-Kitaev-Preskill (GKP) state [47, 57], which is defined as the simultaneous eigenstates of the two stabilizers, i.e.,
$S_{q}=e^{-i2\sqrt{\pi}\hat{p}},\,\,S_{p}=e^{i2\sqrt{\pi}\hat{q}}.$ (17)
The logical operators are correspondingly given by $X_{\scriptscriptstyle L}=\sqrt{S_{p}}$ and $Z_{\scriptscriptstyle L}=\sqrt{S_{q}}$, respectively.
However, the ideal GKP states are unphysical, requiring infinite amount of energy to prepare [47]. Instead, we will consider the approximate GKP states that are defined as (up to normalization):
|$\begin{split}&\ket{\tilde{0}}\propto\sum_{s=-\infty}^{\infty}\int_{-\infty}^{\infty}e^{-\frac{\Delta^{2}}{2}(2s)^{2}\pi}e^{-\frac{1}{2\Delta^{2}}(q-2s\sqrt{\pi})^{2}}\ket{q}\mathrm{d}q,\\ &\ket{\tilde{1}}\propto\sum_{s=-\infty}^{\infty}\int_{-\infty}^{\infty}e^{-\frac{\Delta^{2}}{2}(2s+1)^{2}\pi}e^{-\frac{1}{2\Delta^{2}}(q-(2s+1)\sqrt{\pi})^{2}}\ket{q}\mathrm{d}q.\end{split}$ (18)
Note that the closer the squeezing parameter $\Delta$ is to 0, the higher is the quality of the GKP states [47, 57].
The GKP code is a promising candidate for quantum error correction (QEC) in continuous-variable systems [49], such as trapped-ion mechanical oscillators [56] and superconducting circuits [58]. In addition, GKP states hold the potential for enhanced sensing capabilities, allowing simultaneous detection of displacements in both the position and momentum quadratures of an oscillator [50].
A Oscillator displacements along the momentum and position quadratures
The projectors given in Eq. (14) now become just a displacement operator of the oscillator, which depends on the measurement result of the probe spin, as indicated by:
$\mathcal{P}_{\scriptscriptstyle N=0,\pm}=\frac{D(\theta)\pm e^{i\phi}D(-\theta)}{2},$ (19)
where $D(\theta)=\exp{(\theta a^{\dagger}-\theta^{*}a)}$ represents the displacement operator, with the amplitude $\theta=g_{0}F(\epsilon,t)/2\omega$.
The additional phase $\phi$ is introduced by rotating the probe spin by an angle $\phi$ before the measurement, as shown in the circuit in Fig. 5. The proposed GKP encoding scheme involves repeated execution of this circuit, adaptively setting the angle $\phi$ based on previous measurement results. The angles $\phi$ select the effective measurement basis of the probe spin. This is a phase estimation protocol that incrementally identifies the eigenvalue of $D(2\theta)$ while simultaneously projecting the oscillator into the corresponding eigenstate.
Displacement along position quadrature
Our proposal is realized by carefully tuning the detuning $\epsilon$ and applying a certain number of inversion pulses to the probe spin. For $\epsilon=0$, the filter function simplifies to $F(\epsilon,t)=2(n+2)$ so that $\theta=g_{0}t/\pi$. This leads to conditional displacements exactly along the position quadrature:
$D(\theta)=\exp\Big{[}\mp i\frac{\sqrt{2}g_{0}t}{\pi}\hat{p}\Big{]},\,\,\text{with }\epsilon=0,$ (20)
where $\hat{p}=\frac{i}{\sqrt{2}}(\hat{a}-\hat{a}^{\dagger})$ denotes the momentum operator of the oscillator. According to Ref. [9], this ability to adjust the oscillator’s momentum based on the spin state facilitates a mechanism for entangling two spins by the oscillator.
Displacement along momentum quadrature
Additionally, we also need to realize controlled oscillator-displacements along the momentum quadrature. This is achieved by setting $\epsilon=\omega/(n+1)$ so that $\epsilon t=\pi$, the filter function reads
$F(\epsilon,t)=2+\frac{4}{e^{i\epsilon\tau}-1}=\frac{4}{i\epsilon\tau}+\mathcal{O}(\epsilon\tau),$ (21)
where $\hat{q}=\frac{1}{\sqrt{2}}(\hat{a}+\hat{a}^{\dagger})$ is the position operator of the oscillator. Correspondingly, we have $\theta=ig_{0}t(n+1)/\pi^{2}(n+2)$. Furthermore, neglecting the higher order terms of $\epsilon\tau\sim\mathcal{O}(1/n)$ leads to the following displacement operator of the oscillator:
$D(\theta)=\exp\Big{[}\mp i\frac{2\sqrt{2}g_{0}t}{\pi^{2}}\frac{n+1}{n+2}\hat{q}\Big{]},\,\,\text{with }\epsilon=\frac{\omega}{n+1}.$ (22)
State preparation via phase estimation
[…]
Figure 5: Single round of an adaptive phase estimation protocol that estimates the eigenvalue of oscillator displacement operator, $D(2\theta)=\exp{(2\theta a^{\dagger}-2\theta^{*}a)}$ with amplitude $2\theta=g_{0}F(\epsilon,t)/\omega$. Initially, the ancillary spin is set in the $\ket{+}$ state and subsequently measured in the x-basis after applying a series of periodic inversion pulses, which facilitates the implementation of controlled oscillator displacements. Before measuring the ancilla spin, it undergoes a z-axis rotation by an angle $\phi$. This angle $\phi$ is strategically determined based on the ancilla measurement results of previous phase estimation cycles, with the goal of minimize the uncertainty in the posterior probability distribution estimate of $D(2\theta)$’s eigenvalue. Specifically, when $\phi=0$, this circuit simply executes the projectors given in Eq. (19). Through repeated application of this phase estimation circuit, the oscillator states outside the subspace spanned by the eigenstates of $D(2\theta)$ are gradually suppressed.
Phase estimation has been shown to be effective for preparing such approximate GKP logical states by determining the eigenvalues of stabilizers and logical operators [48]. For the spin-oscillator hybrid system considered here, such eigenvalue determination is realized by adjusting the detuning magnitude $\epsilon$ and the total evolution duration $t$. In this way, we can choose the magnitude by which the oscillator is shifted along the momentum or position quadrature.
For example, to prepare the approximate GKP state |$\ket{\tilde{0}}$, first initialize the oscillator in its ground state with zero photons. It then sets $D(\theta)=S_{q}$ and runs the phase estimation circuit several times to estimate its eigenvalue with high confidence. A subsequent fit to $D(\theta)=Z_{\scriptscriptstyle L}$ allows for similar eigenvalue estimation. Determining the eigenvalues of $S_{q}$ and $Z_{\scriptscriptstyle L}$ projects the oscillator into the desired approximate GKP state. This process can be described as iterative and adaptive application of the circuit in Fig. 5.
GKP encoding and oscillator cooling
The GKP encoding scheme we discuss is similar to the oscillator cooling technique described in Ref. [43], with both methods using the phase estimation circuit shown in Fig. 5 many times. However, the encoding scheme starts from the ground state, as opposed to the cooling method, which applies the circuit to a thermal state and then cools it down to the ground state.
A crucial difference between the two lies in how the circuit is used in the cooling process, where it is implemented with non-zero detuning, so that neither Eq.(16) nor Eq.(18) hold. When these conditions are satisfied, instead of cooling to a zero photon ground state, the oscillator, initially in a thermal state, stabilizes in a state that maintains a finite number of photons [43].
B Experimental feasibility
In contrast to the oscillator cooling scheme developed in Ref. [43], which requires only a weak probe-oscillator coupling, the GKP encoding scheme is much more demanding as it requires a strong coupling strength. In this respect, current experimental capabilities are clearly insufficient to implement the preparation scheme. Moreover, the GKP encoding scheme has been extensively studied in Ref. [48]. Therefore, we choose not to perform detailed numerical simulations to evaluate the practicality of the GKP encoding scheme in this study. Instead, we briefly assess the challenges and discuss potential improvements to existing setups that could facilitate the preparation of GKP states in macroscopic mechanical oscillators.
The time duration for each phase estimation round needs to be of the order of $1/g_{0}$, so that the displacement magnitude can be of the order of $|\theta|\sim O(\sqrt{\pi})$. Given the necessity of around 10 phase estimation cycles to yield an approximate GKP state of acceptable quality [48], the preparation scheme requires an overall time exceeding $\mathcal{O}(10/g_{0})$.
From an experimental point of view, it’s crucial that the coherence time of the oscillator is considerably long compared to the coupling strength $g_{0}$, ideally at least on the order of $100/g_{0}$. Diamond NMOs with a quality factor $Q>10^{6}$ have been demonstrated at sub-Kelvin temperatures and mechanical damping rates of a few kHz [59]. This implies that the coherence time of the vibrational mode must be significantly improved. Note that the coherence time of the NV electronic spin in cryogenic conditions can exceed 1 second by the application of dynamical decoupling [52]. This is far from sufficient, as the probe spin only needs to maintain coherence during each phase estimation round.